# Reasoning with Confidence: Efficient Verification of LLM Reasoning Steps via Uncertainty Heads

### v1

@@ 4.3 Evaluation Settings
UHead Self-anno (Ours) 32K .529 .594 .260 .735 .779 .779 .394 .404 .461 .618 .559
UHead DeepSeek-anno (Ours) 32K .465 .616 .243 .740 .802 .786 .395 .361 .441 .617 .551
5 Results and Analysis
Step-level correctness prediction
results for Qwen-3 are presented in Table 4.3 and for Phi-4 in Table A in Appendix A. We conclude that unsupervised UQ methods in many cases provide valuable signals for detecting reasoning errors. Even lightweight single-sample techniques such as MSP and MaxEntropy demonstrate slight improvements over the random baseline in the majority of datasets, with the exception of calendar planning. Sampling-based methods such as Semantic Entropy usually perform slightly better, but have similar issues on calendar planning.
Small PRMs demonstrate notable improvements over unsupervised UQ methods for mathematical problems and QA, but for OOD planning datasets, they are not better than UQ. Best large PRMs considered in our work, such as Qwen2.5-Math-7B-PRM800k and Qwen2.5-Math-PRM-7B, substantially outperform all unsupervised UQ methods and smaller PRMs on all tasks.
Despite using 750-810 fewer parameters than PRMs, both UHead variants perform on par with, or even surpass the best PRMs. For ID mathematical datasets, UHeads substantially outperform all other PRM baselines except the two strongest PRMs based on Qwen2.5-Math. On MATH, UHead trained on self-supervised annotation is on par with Qwen2.5-Math-PRM-7B. On GSM8K, UHead trained on DeepSeek annotation is on par with Qwen2.5-Math-7B-PRM800k. For ProofNet, UHead falls behind the strongest Qwen-based PRMs, but substantially outperforms all others.
It is not surprising that parameter-heavy PRMs achieve the best results on ID mathematical datasets, as they tend to strongly overfit to this domain. In contrast, UHead, with far fewer parameters, avoids such domain-specific overfitting. This becomes evident on planning tasks, where UHead consistently outperforms the strongest PRMs across all datasets. The best performance is obtained by UHead trained on DeepSeek annotations, though even the self-supervised variant surpasses all PRMs. For QA datasets, the picture is more mixed. On StrategyQA, UHead again emerges as the top step-verification method. However, on ScienceQA, RLHFlow-PRM-DeepSeek makes a clear leap forward, outperforming other techniques, including Qwen-based PRMs. Still, UHead performs on par with Qwen-based PRMs, underscoring its solid performance in this setting.
In summary, UHead lags behind the strongest PRMs on average for ID tasks but outperforms them for OOD tasks. Notably, the average performance of the self-supervised UHead is comparable to that of the externally supervised variant, highlighting its ability to provide an efficient self-supervised solution for reasoning step verification, particularly valuable in OOD settings. Similar results are observed for Phi-4.
Table 3: Offline best-of-$N$ decoding accuracy across datasets (Qwen-3 8B). For datasets with verifiable final answer, we also provide majority voting.
@@ 5 Results and Analysis
UHead Self-anno (Ours) 32K 94.4 97.5 73.6 9.4 6.5 31.0 88.6 97.1 88.5 46.5 62.3
UHead DeepSeek-anno (Ours) 32K 92.7 97.8 76.5 17.2 7.0 26.0 88.6 96.9 89.0 47.1 62.8
Table 4: Online best-of-$N$ decoding accuracy across datasets (Qwen3-8B).
Method # Sample Math (ID) Planning (OOD) QA (OOD) Average
[…]
Math-Shepherd-PRM-7B 440K 72.9 93.7 44.7 5.3 7.7 27.5 65.0 63.4 70.43 33.78 47.53
RLHFlow-PRM-Deepseek-Data 253K 74.0 94.4 48.2 5.1 8.3 28.1 66.8 63.2 72.20 34.30 48.51
RLHFlow-PRM-Mistral-Data 273K 74.1 94.4 46.9 7.5 10.6 29.6 65.0 62.4 71.80 35.02 48.81
Universal-PRM-Qwen2.5-Math-7B 690K 70.4 91.6 43.8 9.1 6.5 28.5 60.4 61.0 68.60 33.10 46.41
Qwen2.5-Math-7B-PRM800k 263K 76.4 91.3 48.0 5.1 11.5 24.5 63.8 63.2 71.90 33.62 47.98
Qwen2.5-Math-PRM-7B 860K 74.0 92.7 55.9 5.0 6.6 26.0 57.6 63.8 74.20 31.80 47.70
[…]
UHead Self-anno (Ours) 32K 74.2 92.9 48.0 9.5 10.1 29.0 69.6 59.8 71.70 35.60 49.14
UHead DeepSeek-anno (Ours) 32K 74.8 93.7 51.5 12.2 13.7 24.6 69.9 58.8 73.33 35.84 49.90
Offline best-of-$N$ results for Qwen-3 are presented in Table 5 (Phi-4 results are in Appendix A, Appendix A). Both UHead types achieve the best offline BoN performance on several benchmarks, including GSM8K, ProofNet, Meeting/Calendar Planning, StrategyQA, and ScienceQA, achieving on par or superior performance to the best PRMs. On MATH and Trip Planning, UHeads rank second-best. Notably, UHead-based BoN selection enables Qwen3-8B to outperform its larger counterpart Qwen3-14B on multiple benchmarks (MATH, GSM8K, ProofNet, Meeting Planning, and ScienceQA).
As in the step-level setting, UHead exhibits strong generalization. While the best PRMs (Universal-PRM-7B, Qwen2.5-MATH-7B-PRM, and Qwen2.5-Math-PRM-7B) reach parity with UHead on in-domain datasets MATH and GSM8K, they often fall behind on OOD tasks, such as planning and StrategyQA. Furthermore, UHead provides stable gains across tasks of varying difficulty, from relatively simple datasets (MATH, GSM8K, ScienceQA) to complex planning benchmarks.
Online best-of-$N$ evaluation results are presented in § 4.3. While PRMs remain highly competitive on ID tasks like MATH and GSM8K, UHead trained on Deepseek-R1 annotations has obtained the runner-up accuracy across all three ID datasets. More impressively, UHeads demonstrate superior generalization to OOD. UHead DeepSeek-anno secures the highest overall and OOD average accuracy, outperforming all PRMs by a considerable margin. By excelling across both ID and OOD tasks with drastically fewer resources, UHeads proves to be a powerful substitute for PRMs as a test-time reasoning guidance tool.
Combining PRM with UHead. PRMs and UHeads capture complementary aspects of reasoning quality: PRMs rely on textual cues from the generated rationale, while UHeads leverage internal uncertainty signals from the LLM itself. To explore whether these two perspectives can be synergistic, we train a logistic regression model that takes both the PRM score and the UHead score as input to predict step-level correctness labels. The model is trained on a random subset of 200 questions from the training set. Table 5 reports the step-level PR-AUC, showing that this combination yields additional improvements. These findings suggest that integrating uncertainty-based and PRM-based signals is a promising direction for future work.
Table 5: Step-level PR-AUC for the experiment combining PRMs with UHead (Qwen-3 8B).
@@
UHead DeepSeek-anno .529 .594 .260
UHead + PRM1 .613 .674 .318
UHead + PRM2 .573 .710 .327
@@
Impact of data quantity and diversity. In practice with budget constraint, it is important to understand the how data quantity and diversity influence UHead performance Figure 2 (top) shows that larger training sets, more questions and sampled reasoning trajectories, consistently improve UHead performance on both ID and OOD tasks. In practice, however, high-quality questions are often scarce, making it difficult to scale by prompt expansion alone. An alternative is to sample multiple reasoning trajectories per question. The bottom row of Figure 2 compares scaling strategies, demonstrating that trajectory sampling is an effective way to boost UHead performance.
@@
To examine the effect of data diversity, we train UHeads on two subsets of 2K questions from the training set: one highly similar and one highly diverse. Questions are embedded using Qwen3-Embedding-8B (qwen3embedding). The similar subset is formed by selecting the 2K nearest neighbors to the median embedding, while the diverse subset is constructed via farthest-first traversal (GONZALEZ1985293). Results in Table 9 show that data diversity benefits UHead performance.
@@
Figure 2: Top row: PR-AUC of UHeads with increasing training set size (x-axis). Bottom row: scaling training data either by adding new unique questions or by sampling additional trajectories. One dataset per domain is shown; other tasks are covered in Appendix A.
Scaling $N$. In Figure 3, we present the best-of-$N$ performance on GSM8k and ScienceQA for different values $1\le N\le 10$. On GSM8k, performance improves with larger $N$ for almost all methods, with both UHead variants consistently achieving the best results across most settings, outperforming the PRM-based baselines. On ScienceQA, accuracy also increases as $N$ grows, but the differences between methods are smaller, and all approaches perform comparably.
@@ 6 Related Work
Process reward models. Research in PRMs has advanced by scaling and refining step-level annotations: from manual labeling (uesato2022solving; lightman2023letsverifystepstep), to MC-based automatic labeling (wang2024mathshepherdverifyreinforcellms; luo2024improvemathematicalreasoninglanguage), and consensus methods combining LLM-as-a-Judge and MC estimation (zhang2025lessonsdevelopingprocessreward; zhao2025genprm). Generative PRMs further extend step-level judgment with long CoT (xiong2025stepwiserstepwisegenerativejudges; zhao2025genprm). In contrast, our approach focuses on cost-efficient step-level verification via LLM internal states.
Uncertainty quantification methods recently have been employed for test-time scaling and improving reasoning performance (mo2024tree; yin-etal-2024-reasoning; zhang-etal-2025-entropy; fu2025deep; yan2025mur; kang2025scalable). However, so far, only weak unsupervised UQ methods have been used. In this work, we propose data-driven UQ techniques that achieve much better performance.
Formal verification has recently been used to verify LLM reasoning steps (zhou2024donttrustverify; hu2025stepproofstepbystepverificationnatural; liu2025safeenhancingmathematicalreasoning; zhou2025stepwiseformalverificationllmbased). However, these methods often require specialized autoformalization data for training and are limited to narrow domains (e.g., math proofs). In contrast, we focus on UQ methods and demonstrate their ability to generalize across domains.
@@ 7 Conclusion
We introduced UHead, a lightweight step-level verifier that reads an LLM’s own internal states to detect incorrect reasoning. UHead can be trained fully self-supervised without human labels, verifiable final answers, or Monte-Carlo rollouts. Across math, planning, and QA, UHead delivers strong in- and out-of-domain results and is competitive with, or better than, far larger PRMs, making it a practical building block for resource-efficient reasoning systems. Beyond replacing PRMs, UHead complements them: combining PRM scores with UHead’s score consistently improves performance, indicating the two signals capture complementary aspects of reasoning quality. This synergy suggests a promising path toward strong hybrid verifiers that marry introspective uncertainty with process rewards. Looking ahead, our findings pave the way for more efficient test-time scaling and self-verification for LLMs in reasoning tasks.
@@
We use publicly available datasets (MATH, GSM8K, ProofNet, ScienceQA, and StrategyQA) which have no data privacy issues. All artifacts we use are under licenses allowing research usage. Human annotations were conducted by the authors of this paper. We do not identify any other ethical risks associated with this study.
Reproducibility Statement
We will fully open-source our trained UHeads, code, prompts, human annotations, and processed datasets to ensure full reproducibility. For all training, evaluation, and sampling, we fix random seeds to 1 or 42 (see specified in scripts). One major challenge of reproducing exact numbers in our tables from scratch is the use of API-based DeepSeek-R1. API-based LLMs are known to be inherently non-deterministic even if fixing prompts and temperature. To address this, we provide all DeepSeek-R1 annotations used in training and evaluation, allowing others to faithfully reproduce our results. If reproducing from scratch, our codebase also guarantees to reproduce similar trends and observations, even if there are slight differences in exact numbers. Code and data: https://anonymous.4open.science/r/uhead-86C1/README.md

### v3

@@ 3 Verifying Reasoning Steps via Probing Internal States of LLMs
4 Experimental Setup
4.1 LLMs for Reasoning, ReProbe Training, Evaluation Datasets, and Baselines
LLMs for reasoning used in our experiments include three state-of-the-art models: Qwen3-8B (Yang et al., 2025) in non-thinking CoT mode, Qwen3-1.7B in native thinking mode, and Phi-4 (Abdin et al., 2025).
Probe training settings. We use Shelmanov et al. (2025)’s framework to train ReProbes. We conducted a vast hyperparameter search; the best training hyperparameters are detailed in § B.2.
Baselines. We benchmark against a broad set of baselines, covering (1) two small 1.5B PRMs; (2) six large state-of-the-art 7-8B PRMs; (3) basic and state-of-the-art UQ methods implemented in the LM-Polygraph framework (Fadeeva et al., 2023; Vashurin et al., 2025). More details about the baselines are presented in § C.1.
Evaluation datasets span three domains: mathematical reasoning (in-domain), planning (OOD), and general knowledge QA (OOD). We select datasets that demand non-trivial reasoning and include reasoning steps that are unambiguously verifiable, enabling reliable evaluation by both LLMs and human annotators. Test dataset details are given in §§ C.2 and C.3.
4.2 Evaluation Settings
We conduct experiments in three settings: (1) step-level correctness assessment; (2) best-of-$N$ TTS; and (3) beam search TTS.
(1) Step-level correctness. For each question in the test set, we generate a CoT trace using the Qwen or Phi models and evaluate the correctness of their reasoning steps with DeepSeek-R1. To ensure accurate judgments, DeepSeek-R1 is provided with the ground-truth answer, reasoning steps, and supporting evidence (see the prompt in § F.2). To validate the reliability of the judge, we evaluate it against (1) the human annotations from a random subset of PRM800k and (2) a manually annotated set of 1000 steps spanning QA, planning, and ProofNet tasks. DeepSeek-R1 achieves 95% acc. on PRM800k and ~90% on other datasets (see Table 13 and Appendix D for details). Based on the judges’ assessments, we compute the PR-AUC evaluation metric.
(2) Best-of-$N$ TTS setting assumes that we generate $N$ reasoning trajectories per problem (N=10N{=}10 for the mathematical and QA datasets, N=5N{=}5 for the planning datasets, temperature=1.0). The evaluation metric in this setting is the accuracy of the final solution. For GSM8K, we use an exact match against the gold-standard final answer. For other datasets, where final answers may be open-ended or structurally complex, we use DeepSeek-R1 to assign binary correctness labels (1 for correct, 0 for incorrect) based on both the problem statement and the reference solution. The grading prompt is provided in § F.2.
(3) Beam search TTS setting assumes that we maintain a beam of $B=5$ partial reasoning trajectories at each step. From each partial trajectory, we sample $N=5$ candidate continuations, resulting in $B\cdot N$ potential candidates at each step. We assign a step-level correctness score to each candidate and retain the top-$B$ expanded trajectories with the highest scores. The correctness of the final solution is evaluated following the BoN setting, using overall accuracy as the metric.
Method # Sample Math (ID) Planning (OOD) QA (OOD) Average
MATH GSM8k ProofNet Trips Calendar StrQA SciQA ID OOD Overall
[…]
Qwen2.5-Math-7B-PRM800k 263K 89.8 80.4 95.2 13.1 45.0 86.7 91.2 88.5 59.0 71.6
Qwen2.5-Math-PRM-7B 860K 88.1 95.4 93.6 13.1 41.5 79.2 83.6 92.4 54.4 70.7
Reasoning Probes (ReProbes)
$\bigstar$ ReProbe, Attn+Logit, Self-anno (Ours) 32K 90.3 95.4 95.1 15.0 41.5 94.8 96.2 93.6 61.9 75.5
$\bigstar$ ReProbe, Attn+Logit, DeepSeek-anno (Ours) 32K 81.8 93.3 79.1 11.8 42.4 90.7 90.8 84.7 58.9 70.0
$\bigstar$ ReProbe, Hidden States, Self-anno (Ours) 32K 84.1 97.3 90.6 15.6 40.5 91.2 92.8 90.7 60.0 73.2
$\bigstar$ ReProbe, Hidden States, DeepSeek-anno (Ours) 32K 86.8 98.8 95.6 13.4 48.0 96.7 96.7 93.7 63.7 76.6
Table 2: Beam search decoding accuracy across datasets (Qwen3-8B).
@@ 4.2 Evaluation Settings
Reasoning Probes (ReProbes)
$\bigstar$ ReProbe, Attn+Logit, Self-anno (Ours) 32K 94.4 97.5 73.6 9.4 6.5 31.0 88.6 97.1 88.5 46.5 62.3
$\bigstar$ ReProbe, Attn+Logit, DeepSeek-anno (Ours) 32K 92.7 97.8 76.5 17.2 7.0 26.0 88.6 96.9 89.0 47.1 62.8
$\bigstar$ ReProbe, Hidden States, Self-anno (Ours) 32K 94.4 96.1 78.5 13.1 6.0 26.0 88.1 95.8 89.7 45.8 62.3
$\bigstar$ ReProbe, Hidden States, DeepSeek-anno (Ours) 32K 92.7 96.1 74.1 15.0 5.5 26.0 88.6 96.9 87.6 46.4 61.9
Table 3: Best-of-NN decoding accuracy across datasets (Qwen3-8B). For datasets with verifiable final answers, we also provide the majority voting baseline.
5 Main Results
Step-level correctness assessment results for Qwen3-8B are presented in Table 1, and for Phi-4 and Qwen3-1.7B in native thinking mode in Tables 6 and 8 correspondingly in Appendix A. Unsupervised UQ methods, while providing valuable signals for detecting reasoning errors, substantially underperform other approaches. Small PRMs show clear gains over unsupervised UQ methods in the mathematical in-domain setting, but yield limited improvements on OOD datasets. The best large PRMs considered in our work, such as Qwen2.5-Math-7B-PRM800k and Qwen2.5-Math-PRM-7B, substantially outperform all unsupervised UQ methods and smaller PRMs on all tasks.
@@ 5 Main Results
It is not surprising that parameter-heavy PRMs achieve slightly better results on ID mathematical datasets, as they tend to strongly overfit to this domain. In contrast, ReProbe, with far fewer parameters, avoids such domain-specific overfitting. This advantage becomes evident in the average OOD PR-AUC, where all ReProbes significantly outperform the best PRMs. For all OOD planning and QA datasets, the best PR-AUC scores are consistently achieved by ReProbes. In summary, ReProbes slightly lag behind the strongest PRMs on ID tasks but outperform them on OOD tasks. Similar results are observed for Phi-4 and for Qwen3-1.7B in the native thinking mode.
Notably, the self-supervised ReProbe achieves comparable average performance to the externally supervised variant, offering an efficient self-supervised solution for reasoning step verification that is particularly effective in OOD settings.
Given the poor performance of unsupervised UQ in identifying incorrect steps, we exclude it from subsequent test time scaling experiments.
Test time scaling. Best-of-$N$ results for Qwen3-8B are presented in Table 3, and supplementary results are in Appendix A. ReProbes achieve the best performance on all datasets except MATH, where ReProbe ranks second best. Notably, ReProbe-based TTS enables Qwen3-8B to outperform its larger sibling – Qwen3-14B on multiple benchmarks: MATH, GSM8K, ProofNet, Meeting Planning, and ScienceQA. As in the step assessment setting, ReProbe exhibits strong generalization. While best PRMs (Universal-PRM-7B, Qwen2.5-MATH-7B-PRM, and Qwen2.5-Math-PRM-7B) reach parity with ReProbe on in-domain datasets MATH and GSM8K, they often fall behind on OOD tasks, such as planning and StrategyQA.
Beam search TTS results with Qwen3-8B are reported in Table 2. Across all in-domain and out-of-domain datasets, ReProbe achieves the strongest performance, outperforming state-of-the-art PRMs.
Overall, ReProbe provides stable gains in test time scaling across tasks of varying difficulty, from relatively simple datasets (MATH, GSM8K, ScienceQA) to complex planning benchmarks.
@@
ReProbe, Attn+Logit, DeepSeek-anno .529 .594 .260
ReProbe + PRM1 .613 .674 .318
ReProbe + PRM2 .573 .710 .327
Table 4: Synergy of PRMs with ReProbes. Step-level PR-AUC, Qwen3-8B.
Method # Sample Math (ID) Planning (OOD) QA (OOD) Average
ReProbe, Attn+Logit, Indiverse 6K .308 .549 .205 .626 .687 .685 .377 .251 .354 .525 .461
ReProbe, Attn+Logit, Diverse 6K .409 .575 .180 .707 .793 .792 .271 .325 .388 .578 .507
Table 5: PR-AUC of ReProbes trained on the diverse vs. homogeneous subsets of 2K questions (with three trajectories sampled per question). Increased data diversity consistently improves ReProbe’s overall performance.
6 Analysis
Synergy of PRMs and ReProbes. PRMs and ReProbes capture complementary aspects of reasoning quality: PRMs rely on their own knowledge and textual cues from the generated rationale, whereas probes exploit internal signals of the LLM that encode model confidence. To explore whether these two perspectives can be synergistic, we train a logistic regression model that takes both the PRM score and the ReProbe score as input to predict step-level correctness labels. The model is trained on a random subset of 200 questions from the training set. Table 4 reports the step-level PR-AUC, showing that this combination yields additional improvements. This finding suggests that integrating confidence-based signal and the PRM external knowledge is a promising direction for future work.
@@ 6 Analysis
Training data quantity and diversity. We investigate how data quantity and diversity influence ReProbe performance. Fig 3 (left) shows that larger training sets, more questions, and sampled reasoning trajectories consistently improve ReProbe performance on both ID and OOD tasks. In practice, however, high-quality questions are often scarce, making it difficult to scale by prompt expansion alone. Alternatively, it is possible to sample multiple reasoning trajectories per question. Fig 3 (right) compares scaling strategies, demonstrating that trajectory sampling is an effective way to boost ReProbe performance.
@@
To study the impact of data diversity, we train ReProbes on two 2K-question subsets of training data: one highly homogeneous and one highly diverse. For similarity assessment, questions are embedded using Qwen3-Embedding-8B (Zhang et al., 2025b). The homogeneous subset is constructed by selecting the 2K nearest neighbors to the median embedding, whereas the diverse subset is obtained via farthest-first traversal (Gonzalez, 1985). Table 5 shows that increased data diversity leads to consistently improved ReProbe performance.
Architecture. We compare three architectures for ReProbe: (1) ReProbe (step-level): leverages a transformer encoder and predicts step-level correctness; (2) Linear Probe (token-level): replacing transformer-based ReProbe with a linear layer to predict token-level labels (all tokens of incorrect step should be labeled incorrect); and (3) ReProbe (token-level): uses a transformer encoder and predicts token-level labels. All variants use the same training setup (hidden states as features, DeepSeek-anno). Table 9 shows that step-level ReProbe outperforms both token-level variants.
@@
Figure 3: Left: PR-AUC of ReProbes with increasing training set size (x-axis). Right: scaling training data either by adding new unique questions or by sampling additional trajectories. Average PR-AUC across all datasets is reported. The results for each individual dataset are in Fig 5, Appendix A.
Figure 4: Best-of-NN accuracy on GSM8k and ScienceQA for different NN values (Qwen3-8B).
Impact of reasoning length on the performance of ReProbes and PRMs is investigated in § A.4. We show that the performance of both PRMs and ReProbes degrades with increasing reasoning length, but only slightly.
Scaling the number of samples $N$. Fig 4 analyzes best-of-$N$ performance on GSM8k and ScienceQA for $1\leq N\leq 10$. As expected, accuracy improves with increasing $N$ across methods on both datasets. On GSM8k, both ReProbe variants consistently outperform PRMs; on ScienceQA all methods exhibit comparable performance.
[…]
@@ 7 Related Work
PRMs. Research in PRMs has advanced by scaling and refining step-level annotations: from manual labeling (Uesato et al., 2022; Lightman et al., 2024), to MC-based automatic labeling (Wang et al., 2024; Luo et al., 2024), and consensus methods combining LLM-as-a-Judge and MC estimation (Zhang et al., 2025c; Zhao et al., 2025). Generative PRMs further extend step-level judgment with long CoT (Xiong et al., 2025; Zhao et al., 2025). Our approach differs from PRMs in that ReProbe is not a language model and operates directly on internal LLM states rather than textual inputs.
Uncertainty quantification methods recently have been employed for test-time scaling and improving reasoning performance (Mo and Xin, 2024; Yin et al., 2024; Zhang et al., 2025a; Fu et al., 2025; Yan et al., 2025; Kang et al., 2025). However, so far, only unsupervised UQ methods have been used as step scorers. We propose a training-based probe and show that it achieves much better performance than unsupervised UQ.
Formal verification has recently been used to verify LLM reasoning steps (Zhou et al., 2024; Hu et al., 2025; Liu et al., 2025; Zhou and Zhang, 2025). However, these methods often require specialized autoformalization data for training and are limited to narrow domains (e.g., math proofs). ReProbes are much more general; our experiments show that they can generalize to OOD.
@@ 8 Conclusion
We introduced ReProbe, a lightweight step-level verifier that reads an LLM’s own internal states to guide reasoning process. ReProbe can be trained fully self-supervised without human labels, verifiable final answers, or Monte-Carlo rollouts. Across math, planning, and QA, ReProbe delivers strong in- and out-of-domain results and is competitive with, or better than, far larger PRMs, making it a practical building block for resource-efficient reasoning systems. Beyond replacing PRMs, ReProbe complements them: combining PRM scores with ReProbe’s score consistently improves performance, indicating the two signals capture complementary aspects of reasoning quality. This synergy suggests a promising path toward strong hybrid verifiers that marry introspective confidence with process rewards. Looking ahead, our findings pave the way for more efficient test-time scaling for LLMs in reasoning tasks.
Limitations
§ 5 shows that ReProbe performance increases with larger training data and benefits from data diversity. Curves in Fig 3 and Fig 5, although show diminishing marginal gains, do not seem to reach the top for tasks like StrategyQA. Therefore, it seems possible to further unleash the potential of ReProbe with further data scaling – sampling more reasoning trajectories per question or adding new questions beyond the PRM800K training set. In this work, we do not involve questions outside the PRM800K training set to establish a fair comparison with PRMs trained on data derived from this set (e.g., Qwen2.5-Math-7B-PRM800K). Due to a limited budget for annotation (annotation of 32K reasoning trajectories with DeepSeek-R1 cost >2000{>}2000 USD), we also do not annotate more reasoning trajectories per question. We leave this promising scaling to future work, including the integration of diverse, high-quality questions outside the math domain.
ReProbes need to be trained on the internal states of the target LLMs they supervise. Therefore, contrary to PRMs, they cannot be directly applied to another LLM since they depend on model-specific internal states. However, ReProbes are highly efficient in parameter size, training data, and training compute, making their training relatively inexpensive. Moreover, once trained, ReProbes can significantly reduce inference costs compared to PRMs. In practice, applications may focus on a small number of widely used models. If ReProbes for these models are shared online (e.g., on HuggingFace), practitioners could readily download and apply them without training from scratch. Future work may also investigate fine-tuning ReProbes to adapt to customized or fine-tuned versions of target LLMs.
@@
We use publicly available datasets (MATH, GSM8K, ProofNet, ScienceQA, and StrategyQA), which have no data privacy issues. All artifacts we use are under licenses allowing research usage. Human annotations were conducted by the authors of this paper. We do not identify any other ethical risks associated with this study.
Reproducibility. We will fully open-source our trained ReProbes, code, prompts, human annotations, and processed datasets to ensure full reproducibility. For all training, evaluation, and sampling, we fix random seeds to 1 or 42 (as specified in the codebase). One major challenge in reproducing the exact numbers in our tables from scratch is the use of API-based DeepSeek-R1. API-based LLMs are known to be inherently non-deterministic even when fixing prompts and temperature. To address this, we provide all DeepSeek-R1 annotations used in training and evaluation, allowing others to faithfully reproduce our results. If reproducing from scratch, our codebase also guaranties the reproduction of similar trends and observations, even if there are slight differences in exact numbers. Code and data: https://anonymous.4open.science/r/ReProbe-86C1/README.md